OpenStudy (hulkart):

A study conducted at a certain college shows that 51% of the school's graduates find a job in their chosen field within a year after graduation. Find the probability that 5 randomly selected graduates all find jobs in their chosen field within a year of graduating. Round to the nearest thousandth if necessary. 0.098 0.068 0.035 2.550 pls I need help

OpenStudy (anonymous):

Wait, i'm doing.

OpenStudy (hulkart):

pls I need help fast?

OpenStudy (anonymous):

I'm getting two answers and hence I'm unable to decide what could be the ultimate answer :(

OpenStudy (anonymous):

Got it

OpenStudy (anonymous):

Do you want it in full procedure or just the answer?

OpenStudy (anonymous):

0.035 is the answer.

OpenStudy (hulkart):

bullseye! But could you wait for at least 10 more questions?

OpenStudy (anonymous):

sure :)

OpenStudy (anonymous):

in the problem. p=51/100 r=5 n=5 and rest else i think you would know how to do. If you dont, then ask me the procedure. i would snap shot it and send u then.

OpenStudy (hulkart):

You are dealt two cards successively (without replacement) from a shuffled deck of 52 playing cards. Find the probability that both cards are black. Express your answer as a simplified fraction. I appreciate your help man...

OpenStudy (anonymous):

Your solution almost done. just wait 30sec

OpenStudy (anonymous):

Thats the way you do it ;)

OpenStudy (anonymous):

Are u there?

OpenStudy (hulkart):

Am here! You are genius

OpenStudy (hulkart):

A charity is conducting a raffle, and it sells one ticket each to 35 women and 21 men. If 5 winners are randomly selected, what is the probability that they are all men? Round to five decimal places. 0.06268 0.00533 0.09537 0.07776

OpenStudy (anonymous):

Dude, its almost done

OpenStudy (hulkart):

You"re the best man!!!!!!!!!!

OpenStudy (anonymous):

Mean time you also solve for the answer. I'm also doing, i'm making some silly mistake in calculation. If u got doubt to solve, then ask me. okay?

OpenStudy (hulkart):

just solve it pls I'm really bad at math....I have no room for mistakes, It"s really important please man

OpenStudy (anonymous):

0.0053272199498615 ENJOY :D

OpenStudy (anonymous):

Feeling good wit the number ? ;)

OpenStudy (anonymous):

It was said round to five decimal places, show them that you rounded to fifteen decimal places which is going to get them feel GROUNDED :P

OpenStudy (hulkart):

Of course lol I"ve got a lot more pls be patient, add me so you can help me tomorrow A tourist in France wants to visit 7 different cities. If the route is randomly selected, what is the probability they will visit the cities in alphabetical order? 5040...

OpenStudy (anonymous):

There are 7! Ways to arrange the routes. 7! = 7×6×5× 4×3× 2×1 = 5040 Because there is only one arrange that will be in alphabetical order, the probability is 1/5040. Done. Next please :-P ?

OpenStudy (hulkart):

"It will definitely get dark tonight." 1 0.67 0.5 0.30 you were correct!

OpenStudy (anonymous):

what is the meaning of that?

OpenStudy (anonymous):

i see no mathematics in that.

OpenStudy (hulkart):

(this should first, that was a mistake)Express the indicated degree of likelihood as a probability value.

OpenStudy (anonymous):

Answer is 1 :)

OpenStudy (anonymous):

explanation: he purely said it is going to get dark, hence 1, if he wud have mentioned any percentage of darkness, then the value might have changed. So until then, its surely 1 and only ONE :P

OpenStudy (hulkart):

What is the likelihood that a fair coin will land heads or tails? 0.25 0.50 1 0 *hehe great! just tell me whenever you want togo offline or break till tomorrow, ok? If I pass this test I will let you know...I hope I do

OpenStudy (anonymous):

0.5 Buddy, let's finish it off :) I'm with you til the end ;) I will surely intimate you if i'm about to leave :) Happy ?

OpenStudy (anonymous):

Come, on, Let's solve. Hurry up, you have to finish it off and rank top in class :P

OpenStudy (hulkart):

Thank you soo much!! One hundred employees of a company are asked how they get to work and whether they work full time or part time. The table below shows the results. If one of the 100 employees is randomly selected, find the probability that the person drives alone or bicycles to work. Full time Part time Public Transportation 9 8 Bicycle 3 4 Drive alone 32 28 Carpool 9 7 0.60 0.35 0.67 0.63

OpenStudy (anonymous):

You want procedure for that or just the answer?

OpenStudy (hulkart):

just answers PlEASE!!!!

OpenStudy (anonymous):

Never mind, i will give u full procedure :P Here you go ;) How many people drive alone? 32 + 28 = 60 How many people cycle? 3 + 4 = 7 So the number we want is 60 + 7 = 67. So what is the probability? 67 / 100 = 0.67.

OpenStudy (hulkart):

wow that"s just amazing! A circle is divided into 9 uneven "pie" shaped pieces. Each section is labeled 1, 2, ... , 9. The circle is then spun and the number pointing north recorded. Estimate the probability that the next number pointing north will be between 2 and 7 inclusive. 4 3 2 1 9 8 5 6 7 4 3 2 1 5 8 6 5 4

OpenStudy (anonymous):

Since they are uneven pieces we find experimental probability. Section : Frequency 3: 2 4: 3 5: 3 6: 2 7: 1 total frequency between 2-7 inclusive: 13 total frequency: 18 probability estimate = 13/18 Done, as simple as that :P what next hero ? ;)

OpenStudy (hulkart):

20 more question to go! so you see we'll have...I mean 'you', hehe Eight basketball players are to be selected to play in a special game. The players will be selected from a list of 27 players. If the players are selected randomly, what is the probability that the 8 tallest players will be selected assuming none of the players are the exact same height?

OpenStudy (hulkart):

...we'll have to to take a break

OpenStudy (anonymous):

Uhuh, sure then :) The moment you say. we would take a break then :P You say when, i would follow then ;)

OpenStudy (anonymous):

= 8C8 / 27C8 = 1 / 2220075 answer for that basketball problem ;)

OpenStudy (hulkart):

not yet....if you can wait till 9pm( american time)

OpenStudy (anonymous):

just another 1 hour, is it ? Now its going to b 8 there. Right?

OpenStudy (hulkart):

yeah like 7:59Of the 84 people who answered "yes" to a question, 14 were male. Of the 73 people that answered "no" to the question, 12 were male. If one person is selected at random from the group, what is the probability that the person answered "yes" or was male? Round to the nearest thousandth. 0.701 0.166 0.167 0.611 keep up the good work, man!

OpenStudy (anonymous):

(84 yes + 12 no male)/(84+73) Answer is : 0.6114649682 Next problem pLeazzz :P

OpenStudy (hulkart):

Nice! There are 6 members on a board of directors. If they must elect a chairperson, a secretary, and a treasurer, how many different slates of candidates are possible? 720 120 20 216

OpenStudy (anonymous):

There are 6 choices for chairperson, 5 for secretary (chairperson is already taken) and 4 for treasurer(chairperson and secretary already picked. So, the maximum amount of slates is 6 x 5 x 4 = 120 Hehe, nice problem ! Release the Beast (Hehe, i mean nexxt problem :P )

OpenStudy (hulkart):

WOW you're Awesome The organizer of a television show must select 5 people to participate in the show. The participants will be selected from a list of 30 people who have written in to the show. If the participants are selected randomly and none of the 30 people are the same age, what is the probability that the 5 youngest people will be selected?

OpenStudy (anonymous):

can you get me options? Am little small doubtful :P

OpenStudy (anonymous):

denominator would be 142506 I need to see answer's whether the numerator is 5 or 1.

OpenStudy (hulkart):

1/2 3/5 5/3 2/3

OpenStudy (anonymous):

are u sure they are the options? or did u modify them to fractional form?

OpenStudy (hulkart):

they are.

OpenStudy (hulkart):

let's just proceed...

OpenStudy (anonymous):

Dude, am damn sure. The answer is 1/142506 No doubt. the options are wrong. Trust me ! Okay, lets proceed. But count on my answer. Am definetly correct the answer is 1/30C5 which is 1/142506

OpenStudy (hulkart):

you are goddamn right. Two events, A and B, are disjoint. What is the probability that both A and B will occur at the same time? 0 1 0.5 And I trust you!

OpenStudy (anonymous):

If two events are disjoint,then the probability of them both occurring at the same time is 0. Disjoint: P(A and B) = 0

OpenStudy (anonymous):

next question !

OpenStudy (hulkart):

haha, You are the real MVP! Event A has a 0.3 probability of occuring and event B has a 0.4 probability of occuring. A and B are independent events. What is the probability that neither A or B occurs? 0.42 0.10 0.70 0.12 honestly, how do your classmates catch up!?

OpenStudy (anonymous):

P(~A&~B) = 1 - 0.58 = 0.42 Done in just a charm :P Well, you are right, they never catch up, cuz i'm the last grade in the class (Haha, for real :D)

OpenStudy (hulkart):

A local factory uses a manufacturing process in which 70% of the final products meet quality standards and 30% are found to be defective. If three of the products are randomly selected, find the probability that all of them are acceptable. 0.027 0.343 0.429 2.1 so do you mean they"re smarter!?

OpenStudy (anonymous):

0.7^3 = 0.3430 May b, they are. I pretend to be the Dumb guy always :I

OpenStudy (hulkart):

How many ways can 6 people be chosen and arranged in a straight line if there are 8 people to choose from? 20,160 720 40,320 48 In the U.S. it"s different...

OpenStudy (anonymous):

8!/(8 - 6)! = 8!/2! = (8 x 7 x 6 x 5 x 4 x 3 x 2 x 1)/(2 x 1) = 8 x 7 x 6 x 5 x 4 x 3 = 20160 Well, it's atleast better then cuz it's just different out there. But over here in INDIA it's way toooooo.... different. It's Really weird ! You would not want to know it :P

OpenStudy (hulkart):

In one town, 79% of adults have health insurance. What is the probability that 4 adults selected at random from the town all have health insurance? Round to the nearest thousandth if necessary. 0.051 0.390 3.164 0.79 I read slumdog milloniare(a book) it was awesome, I watch bigbang theory and raj is funny...tell me more about India

OpenStudy (anonymous):

Haha, getting interested about india ? :D Well, there's a lot to say (THE FUNNY AS WELL AS STUNNY ;) :P ) So, for you to better understand, just type "it happens only in india" to see the funny side of country and "greatness of india" , "incredible india", "vedas and their values", "traditional dances in india", "sanskrit language histroy", etc, etc :P Answer for above health insurance question is: (0.79)^4 = 0.38950081

OpenStudy (hulkart):

How many ways can an IRS auditor select 5 of 10 tax returns for an audit? 100,000 252 120 30,240 hehe...so what do you want to major in college?

OpenStudy (anonymous):

I'm already holding a Bachelor of Technology degree in Aeronautical Sciences (SEM-2) . But I'm quite not happy with my self about it. So planning to write an All india test for scoring a seat in renowned universities called IIT's :P 10C5 = (10*9*8*7*6)/(1*2*3*4*5) = 252 ways

OpenStudy (hulkart):

WHAT? 0.O America needs people like you. A man flips a coin twice, and each time gets either heads (H) or tails (T). Create the sample space of possible outcomes. HH TT HT HT HH HT TH TT HH HT TT HT TH you are so humble!! You must have tons of friends....

OpenStudy (anonymous):

Well. its to the weirdest people, things always flip off instead of dicing off :P It's like, i get either the worst, or only the best (hehe, probability 1/2):P Hence you know, conclusion: Believe me, i ain't got even one friend. Though lots of people are around me (-_-) i just dont like to call anybody "a friend of mine." Cuz i find most of the people posing (indeed true expressions) but only for an amount of time, later they change and forget people slowly (its part of growing i accept). But i feel low that they do not care anymore. So u know, i just stay calm and continue my way of life hence ! Answer: HH HT TH TT

OpenStudy (hulkart):

4 questions left! A firm uses trend projection and seasonal factors to simulate sales for a given time period. It assigns "0" if sales fall, "1" if sales are steady, "2" if sales rise moderately, and "3" if sales rise a lot. The simulator generates the following output. 0 1 0 2 2 0 0 1 2 3 2 0 2 0 2 2 1 2 3 1 2 2 2 0 3 0 0 2 1 2 1 Estimate the probability that sales will rise moderately. 0.633 0.452 0.419 0.312 right now I have no friends too, except on social medias....so, I kinda feel you bro just like you've got your mathematics ability to cheer you, so have I got my talent...I draw

OpenStudy (anonymous):

Dude, in tht question is it rise moderately or atleast rise moderately ? Cuz the both make a lot of difference to the answer !

OpenStudy (hulkart):

rise moderately. after I recieve your I'll have to hit the sack , pls don't dissapoint tomorrow...my college depends so much on this tests...The truth is that you just gave me hope, that I can do this....or we...I mean 'you'! lol :)

OpenStudy (hulkart):

...recieve your response

OpenStudy (anonymous):

0.4193548387096774 answer ! Am sorry i was late to response, actually i typed a big response but suddenly copied it and forgetting it i copied off another text. so lost the response -_- am sorry !

OpenStudy (anonymous):

well, sure, we wud do your works whenever you want :P

OpenStudy (hulkart):

:) *thumbs up*

OpenStudy (anonymous):

So, thats it huh? All done? Anything still ? :) I did a whole night out solving your problems :P Now i will sleep whole day hence , hehe !

OpenStudy (anonymous):

I see no sign of you over here, are u logged out or what?

OpenStudy (hulkart):

ok...I was trying to find my history so we can continue here Which event below is independent of event A? A: You get to work on time. There are no accidents on the road. It's snowing outside. You remember to pick up milk on the way home. Your alarm goes off. *It may take two weeks to finish these tests, are you really willing to help me?? I know It's too much...